OpenStudy (anonymous):

an errand boy went to the bank to deposit some bills for his employer. Some if the bills were one-dollar bills and the rest were five-dollar bills. The total value of the bills was 86$ and the number of bills was 38. Find the number of each kind of bill. system of equations. Please help if you can thanks!

OpenStudy (anonymous):

ones + fives = 38 (1)(ones) + (5)(fives) = 86 If you don't like "ones" and "fives", you can use "x" and "y" if you like. You can solve this now by either elimination or substitution.

OpenStudy (anonymous):

Are you able to solve the above system I identified for you, @maddie227 ?

OpenStudy (anonymous):

um... yeah it helped a bit but still not coming up with an answer.

OpenStudy (anonymous):

Try using "x" and "y". Let "x" be the number of "one dollar bills" and let "y" be the number of "five dollar bills". x + y = 38 x + 5y = 86 x = 38 - y x + 5y = 86 rearranging 1st equation (38 - y) + 5y = 86 substituting 1st into 2nd 38 + (-y + 5y) = 86 associative law 38 + 4y = 86 simplify 4y = 48 y = 12 So, x = 26

OpenStudy (anonymous):

And that's all there is, and now you have the # of one and five dollar bills (x and y). All good now, @maddie227 ?

OpenStudy (anonymous):

thanks! @tcarroll010

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! You're welcome, @maddie227